A Motorcycle Catches a Car A motorcycle is follow

I need help with both parts
Answer #1

PART A)

initial relative speed, u= 23-23 =0

relative acceleration, a = 5 m/s^2

relative distance, s = 58 m

using second equation of motion,

s = ut + 0.5at^2

58 =0+0.5*5t^2

t^2 = 2*58/5 =23.2

t = 4.82 s

Hence t2 - t1 = 4.82 s

part B)

using second equation of mottion,

s = ut +0.5at^2

   = 23*4.82 + 58

     = 169 m